Daily Reports

Because of COVID-19, only IRAM staff is allowed to observe on-site until new notice.

Back to the NIKA2 Pool page

Nov 23rd, Tuesday

We successfully arrived at the telescope with the first winter transport of the season.

Afternoon shift (16:00-21:00; Ioannis + Stefano): we try to focus but the sky is highly variable. We give up and stop when the peak-to-peak variations in the timeline of pointings reach 400 Jy/beam (four hundred). We're now immersed in the clouds. Temperature -9 Celsius; humidity 100%, wind ~10 m/s.

Today's journey to the stars:
attachment:IMG_3034_mod_sm.JPG attachment:IMG_3052_mod_sm.JPG attachment:IMG_3071_sm.JPG

First half of the night shift (21:00-03:00; Stefano and Juan): fog, snow and wind. The telescope is parked. The de-icing is on. At roughly midnight (CET) the wind calms down and it is not snowing anymore. It seems that the fog might also disappear soon. At 00:22 CET the sky is clear and the telescope moves out of the parking position. We now need to wait a bit until the primary surface relaxes after the de-icing. Frederic goes up on the platform to remove some snow and ice from the subreflector. After the usual pointing/focus, we switch to 010-21 and we observe 3 repetitions of NGC1333-C2, thus reaching a total of 17 repetitions and 5h40m in total on this source (total time requested 7h per target). Tau(225)~0.20-0.30. Calib_1scan on Uranus (20211124s21)

Nov 24th, Wednesday

03:00 - 09:00, Longji: The shift started with the observation on 010-21. After 1-hour observations, a new focus and a skydip (20211124s40), I switched to N2CLS(192-16) with tau~0.27. The weather degraded quickly in less than 1 hour and I stopped the observation. I switched to 045-21 from 6:30am CET. A new focus was made at the end of the shift but we have elliptical beam at 1mm even at the best focus returned.

09:00 - 15:00, Ioannis & Carsten: tau~0.37, +/-5Jy/b: 2 scans on NGC3938 (IMEGIN), Switching to 086-21 (ARSENAL, 1xrepetition). Entering into clouds and becoming much more instable. Switched to EMIR backup 015-21. Mars peak temperature varies by >50%. Stop.

15:00 - 21:00, Andrew (remote): The telescope has remained in fog throughout this shift. I didn't enter a single command into Pako. Better luck to the next shifts... :*-(

21:00 - 03:00, Stefano & Laurence: Bad weather and telescope parked throughout the shift.

All we could observe today...

Nov 25th, Thursday

03:00 - 09:00, Alex waking up, bad weather telescope in park in the fog, waiting for clear skies. 07:20, Sky is clearing up, no instruction on what to observe (wrong email address) so doing a focus on 1226+023 to proceed on 192-6... tau at 0.17. Took some time to get a proper focus (mostly in my brain)

09:00 - 15:00, Carsten & Ioannis. tau~0.25-0.3 stable. Clear sky, some highy ice clouds. Doing 084-21. Source at ~30deg from the sun. +-5Jy/b at 2mm. calib_1scan #75 on Mars. Skydip #76. Continuing on 084-21 until thick clouds rise leading to +-20Jy/b and more of instability. Stopped.

15:00 - 21:00, Andrew (remote). After 90 mins, the fog cleared and turned into some excellent conditions. Two full sweeps of the GASTON GP field with tau ~0.15 before it got too low - the GP field is almost complete now (~1 hour remaining or so). Skydip [20211125s129], and calib_1scan [20211125s138] on Uranus before moving to 007-21. The weather deteriorated rapidly during the second repeat on 4149, which is likely to have some junk data as a result.

Our favorite telescope observing mapping the Galactic plane for GASTON...
(Download and watch it bigger on your laptop)

21:00-03:00 Stefano + Juan + Philippe + Michael + Quang. The telescope is stopped. Around 23 UT Stefano suggests all others to go to bed and have some good rest. The sky is highly variable. We have also some consistent snow fall. At ~1 a.m. UT, it does not snow anymore and the de-icing is then switched off. Tau is still quite unstable and there's still cloud coverage. Let's see how it develops in the next hour or so... More snow!

Nov 26th, Friday

03:00 - 09:00 Guilaine: Start at around 4h15. Pointing/Focus. KIDs tuning not good. Atmosphere very unstable. Keep on trying. Around 05h20 real improvement. Tau around 0.3 but not very stable. Good focus. Do one repeat on A1413 (086-21, switching to 48Hz). 6h25: tau improved much faster than predictions, go to N2CLS (cosmos field, 2x(2x27) min) [acquisition back to 24 Hz]. Focus at 7h30 and 8h40. Last pointing at 8:40. Launch one more repeat on N2CLS (tau = 0.11, wow!).

09:00 - 15:00 Ioannis: After pointing and focus on 1253-055, 2 repetitions for project 084-21. Then pointing and focus on Mars, as well as a calib_1scan [20211126s94] and a skydip [20211126s95]. Afterwards, two more repetitions for 084-21 and the project is completed. Very good weather throughout the shift with tau ~ 0.15.

15:00 - 21:00 Andrew (remote). Pointing and focus on 1741-038, 1mm beams are slightly elliptical. With an excellent tau of ~0.15, I completed the GASTON Galactic Plane survey field after just over 4 years since the first scans (nice!). Three repetitions on B335 for project 035-21 with tau increasing to ~0.25 for the last scans. After a skydip (20211126s156) and calib_1scan on NGC7027 (20211126s157), I moved on to 007-21 and 4149.

21:00 - 03:00. Laurence, Emmanuel & Stefano. The sky keeps good/decent. We observe one repetition on Cluster PSZ1G080 (199-16), then calibration scans on Uranus (calib_1scan is 20211126s193; beammap is s195); then we switch to 007-21 and observe sources 4149 (1 repetition before it gets too high in Elevation) and source 1390 (7 repetitions - all repetitions of this source have been completed!).

Today's sunset :-)
attachment:IMG_5065_mod_sm.JPG attachment:IMG_5619_mod_sm.JPG

...and starry night!

Nov 27th, Saturday

03:00 - 09:00 Ankur: All reps ( 18/18) on source 1390 of project 007-21 completed at beginning of shift. Tau was around 0.23 and did a skydip (s27) and calib_1scan (s30) on Uranus. As tau dropped to ~0.15, started with 085-21 project and did 4 reps on source eFEDS3M. Very stable conditions +/- 4Jy/Beam throughout. Also did a beammap on source 1226+023 (s54). Tau ~ 0.21 towards end of shift, continues to be stable condition.

09:00 - 15:00. Carsten (on site): calib_1scan on Mars (#77), skydip (#78), E03-20 with 20"/s (#80-85) as the source was >76.5deg. tau~0.25. +/-6-8Jy/b @ 2mm. Starting with #87 back to 40"/s at elevations < 76deg. tau slowly improving to 0.20. Staying rather stable with less than +/-8Jy/b variability at 2mm. 12:30: Clouds coming. tau~0.18, but becoming much more instable: +/-30Jy/b. 12:50: thick local clouds. stopped.

15:00 - 21:00. Andrew (remote). No observations due to poor conditions.

21:00 - 03:00. Stefano (+ Philippe + Michael + Quang). Wind and snow. No observations.

Nov 28th, Sunday

03:00 - 09:00 Ankur: Telescope started at around 06:45. Tau ~ 0.11 and stable. +/- 4 Jy/Beam on 2mm. Skydip (s15) done. Started with 085-21 but cancelled and telescope stopped due to high wind.

09:00 - 15:00 Carsten: Started again at 09:35. Wind much lower. tau~0.11, +/-8Jy/beam at 2mm. Mars: calib_1scan s26. Skydip (s27). Beam map (s28). Switching to E03-20. Pointing jump by 4" confirmed. Weather becoming very stable: better than +/-2Jy/b, then slowly degrading again. Mars calib_1scan: s94.

15:00 - 21:00 Andrew (remote): Began observations for 035-21 after pointing and focus, with tau~0.18. Completed B335, and observed L694 until moving onto 007-21 when source 4149 was high enough, and stayed there for the rest of the shift. Weather slowly degraded from ~0.18 at the beginning to ~0.27 by the end. 1mm beams slightly elliptical early on, but were pretty much perfect for 007-21.

21:00 - 03:00 Marco, Antonio and Stefano: We let the 007-21 scans to finish for 15 minutes and then we switch to 199-16. We start with opacity around 0.3 but stable (+/-1 Jy/beam). We perform 1,5 repetitions of PSZ1G080 [2/5, 4 scans/each] but then tau becomes too high [> 0.35]. The sky still has very good stability but the opacity is too high for 199-16 (~0.35 at this point). It is too high also for 007-21. The sources of 010-21 are too high in elevation. We are left with the time fillers. We opt for 045-21 because it does not need any reboot of the DAQ (086-21 needs 47.6 Hz sampling). At 00:00 UT we have to park the telescope because we're in the fog/cloud and it could well freeze onto the main dish.

Today's wonderful clear skies
attachment:IMG_6917_sm.JPG attachment:IMG_7596_mod_sm.JPG

Nov 29th, Monday

03:00 - 09:00 Arkur: Telescope stowed due to thick clouds and some snow fall. De-icing switched-on.

09:00 - 15:00 Ioannis, Carsten: Cleared-up. We do 2 repetitions for project E03-20. Afterwards tau225 increased (>0.35) and atmospheric instabilities increased. We change to ARSENAL (project 086-21) and do two repetitions on A2163. The weather improves and stability gets better. We return to project E03-20 for another two repetitions. Some of the E03-20 scans might not be usable because the "auto-tune" option of the acquisition was deactivated by mistake close to the end of the shift. This was noticed and corrected a bit later (s82 was the last one affected).

15:00 - 21:00 Andrew (remote). Observed 035-21 when I took over. Tau was high and variable (~0.3-4), but the 2mm timelines were stable, with +/- 2 Jy/beam. One full script on L694, and then 1 further scan before cancelling the remaining 3 when the weather really degraded. I switched to 045-21, observing the full block of 5 observations at GP052_3. Moved to 007-21 when the weather improved again, and after some focus issues around sunset. 1 script on 869, and two on 4149.

21:00-03:00 Philippe, Michael & Quang (+ Stefano). Observations of 122-16 & 018-21 [8 maps in GASTON Taurus sources, 4 maps in 018-21 (Persus)]. Tau~0.2 throughout the whole shift. skydip (20211129s154), calib_1scan (20211129s161), Beammap(s162).

Spot the planets and constellations:
attachment:IMG_10940_sm.JPG attachment:end2sm.jpg

Nov 30th, Tuesday

03:00 - 09:00 Ankur: Tau ~ 0.2 at 03h and was stable (+/- 2Jy/Beam on 2mm) throughout. Skydip (s13) , Calib_1 scan (s16) and beammap (s17) on Uranus done. Completed 3 reps on project 085-021. Completed 2 reps on source 4149 of project 007-21. Finished with 1 rep on project E03-20.

09:00 - 10:30 Ioannis: Tau225 shows small fluctuations around 0.15. E03-20 repetitions not completed due to high-elevation. Going to Mars for focus, calib_1scan [20211130s73] and beammap [20211130s74]. Then skydip [20211130s75] and back to E03-20 for one last repetition (scan stopped because of the start of weekly maintenance, project almost completed).

Dec 1st, Wednesday

03:40 - 06:00 Manuel, Miguel, Angel: NIKA2 pointing session under technical time (T07). Superb conditions (tau255 < 0.15; high stability). Focus/first pointing on W3OH. A total of 24 pointing scans were obtained (20211201s46 to s69; s68 aborted). Data file sent to Manuel Castillo for P* values fitting. Data stored in ~/mrt-lx3/vis/nikas-21/observationData/imbfits/, however it had to be done from the nikaw-21 account (not yet configured).

DailyReportsNika2Pool231121 (last edited 2022-03-23 10:50:12 by NikaBolometer)